OpenStudy (a.somera32):

how do i find the exact value of the trigonometric function -> cos 9π/4

OpenStudy (zzr0ck3r):

\(\cos(\dfrac{9\pi}{4})=\cos(\dfrac{9\pi}{4}-2\pi)=\cos(\dfrac{\pi}{4})=\ ?\)

OpenStudy (welshfella):



TheSmartOne (thesmartone):

just to elaborate: pi/4 = pi/4 * 180/pi = 180/4 = 45 degrees we multiply bu 180/pi to change from radians to degrees since 1 pi radians = 180 degrees and you know that 45-45-90 triangle has side ratios of 1, 1, sqrt(2) and you should know the following: sin = opposite/hypotenuse cos = adjacent/hypotenuse tan = sin/cos = opposite/adjacent sec = 1/cos csc = 1/sin cot = 1/tan = cos/sin

OpenStudy (welshfella):

Also the reason why 2pi was subtracted from 9pi/4 was because the cosine is a periodic function where the same value recurs every 2pi radians.

OpenStudy (sshayer):

\[\cos \frac{ 9 \pi }{ 4 }=\cos \left( 2 \pi+\frac{ \pi }{ 4 } \right)=\cos \frac{ \pi }{ 4 }=?\]